Use the normal distribution of SAT critical reading scores for which the mean is 514 and the standard deviation is 123. Assume the variable x is normally distributed.


 (a) What percent of the SAT verbal scores are less than 600?

 (b) If 1000 SAT verbal scores are randomly selected, about how many would you expect to be greater than 550?


 (a) Approximately _______ % of the SAT verbal scores are less than 600. 

(Round to two decimal places as needed)
Answer #1

(a)

= 514

= 123

To find P(X<600):

Z = (600 - 514)/123 = 0.6992

Table of Area Under Standard Normal Curve gives area = 0.2580

So,

P(X<600)= 0.5 + 0.2570= 0.7580 = 75.80%

So,

Answer is:

75.80 %

(b)

= 514

= 123

n = 1000

SE = /

= 123/= 3.8896

To find P(>550):
Z = (550 - 514)/3.8896 = 9.2555

Table of Area Under Standard Normal Curve gives area = 0.5 nearly

So,

P(>550)= 0.5 - 0.5 nearly = 0 nearly = 0.00 %

So,

Answer is:

0.00 %